# Zipper-LoRA: Dynamic Parameter Decoupling for Speech-LLM based Multilingual Speech Recognition

### 18 Mar 2026 v1

@@ III-B Representative PEFT Adaptations of Speech-LLM
where $A\in\mathbb{R}^{r\times d_{in}}$ and $B\in\mathbb{R}^{d_{out}\times r}$ are shared low-rank factors (down-/up-projection), and $\alpha$ is a scaling hyper-parameter. Since both $A$ and $B$ are fully shared across languages, the same low-rank subspace is used to adapt $W_{0}$. While this promotes maximum cross-lingual knowledge transfer, it is highly susceptible to negative inter-lingual interference. Under imbalanced multilingual training, the shared subspace tends to be dominated by high-resource languages, which can suppress learning signals for low-resource languages [46].
@@
FlyLoRA: A more recent advancement, FlyLoRA [57], introduces a rank-wise mixture-of-experts (MoE) mechanism into the adaptation process. As shown in Fig. 3(c), it treats each rank component as an individual expert and dynamically activates a sparse subset of ranks via an implicit routing mechanism. Specifically, it employs a sparse and frozen down-projection matrix $A$ to compute routing scores $\mathbf{y}=A\mathbf{x}+\mathbf{d}$, where $\mathbf{d}\in\mathbb{R}^{r}$ is a learnable bias term for load balancing. Based on these scores, a top-$k$ operator selects the active indices $\mathcal{I}_{\text{top}k}\subseteq\{1,\dots,r\}$. The adapter weight is then constructed by aggregating the top-$k$ active components:
@@ IV Proposed Zipper-LoRA
To harmonize these objectives, we propose Zipper-LoRA, a family of PEFT modules designed to explicitly allocate and integrate shared and language-specific rank components in a language-aware manner. As illustrated in Fig. 4, all Zipper-LoRA variants leverage a globally shared down-projection matrix $A$, ensuring common acoustic-linguistic feature extraction. The fundamental distinction between these variants lies in the rank-wise construction of the up-projection matrix $B$. Intuitively, Zipper-LoRA “zips” together shared and language-specific columns of $B$ to form an effective rank-$r$ adapter for each language. This integration can be realized through either a static allocation or a dynamic rank-wise composition.
To provide a comprehensive understanding of the proposed framework, in the following sections, we first establish the definition of the shared down-projection matrix $A$ and the rank-wise construction of $B$. We then present the proposed Zipper-LoRA family, ranging from the static language-wise hard split to dynamic binary selection and soft mixing mechanisms. This is followed by the LID-aware contextual routing that manages these rank-wise compositions, concluding with the specialized training strategies tailored for optimizing Zipper-LoRA, such as segment-wise encoding and parameter initialization.
IV-A Definition: Shared $A$ and Rank-wise Construction of $B$
@@ V-B Configurations
We use Whisper Large-v3 [37] as the speech encoder and Qwen3-1.7B [50] as the decoder-only language model. The modality projector adopts a gated projection module that performs temporal downsampling by a factor of $4$. Specifically, it applies two 1D convolution layers (kernel size $4$, stride $4$) to produce a gate branch and an up-projection branch, fuses them via a SiLU-gated interaction [11], and then applies a linear transformation followed by a residual connection with layer normalization. A final linear projection outputs the speech-conditioned embedding sequence. Combined with the initial $2\times$ downsampling in the speech encoder, this results in an overall $8\times$ temporal reduction, yielding a highly compressed representation at $12.5$ Hz that is fed into the language model.
For parameter-efficient adaptation, we insert low-rank adaptation modules into all linear layers on both the encoder and language-model sides, including the attention projections and feed-forward network layers. We set the rank to $32$ and the scaling factor to $64$ for all methods. For FlyLoRA [57], we activate $8$ rank components. For the chunked encoder regime, we randomly sample chunk durations from $\{1,2,4,8\}$ seconds during training and use a fixed chunk duration of $8$ seconds during decoding; no overlap is used. The non-chunked regime encodes each utterance with full context. All experiments are trained on $4$ NVIDIA A800 GPUs (80 GB) with DeepSpeed ZeRO stage $2$ [38], using a learning rate of $2\times 10^{-5}$, a warmup ratio of $10\%$, and a cosine learning-rate schedule. To improve memory efficiency, we pack training samples by constraining the total sequence length to at most 2048 tokens, where the packed length includes the downsampled speech tokens, the text tokens, and reserved prompt tokens; packing is performed within the same language. After packing, we use a batch size of $16$ for chunked training and a batch size of $10$ for non-chunked training.

### 19 Mar 2026 v2

@@ III-B Representative PEFT Adaptations of Speech-LLM
where $A\in\mathbb{R}^{r\times d_{in}}$ and $B\in\mathbb{R}^{d_{out}\times r}$ are shared low-rank factors (down-/up-projection), and $\alpha$ is a scaling hyper-parameter. Since both AA and $B$ are fully shared across languages, the same low-rank subspace is used to adapt $W_{0}$. While this promotes maximum cross-lingual knowledge transfer, it is highly susceptible to negative inter-lingual interference. Under imbalanced multilingual training, the shared subspace tends to be dominated by high-resource languages, which can suppress learning signals for low-resource languages [46].
@@
FlyLoRA: A more recent advancement, FlyLoRA [57], introduces a rank-wise mixture-of-experts (MoE) mechanism into the adaptation process. As shown in Fig. 3(c), it treats each rank component as an individual expert and dynamically activates a sparse subset of ranks via an implicit routing mechanism. Specifically, it employs a sparse and frozen down-projection matrix AA to compute routing scores $\mathbf{y}=A\mathbf{x}+\mathbf{d}$, where $\mathbf{d}\in\mathbb{R}^{r}$ is a learnable bias term for load balancing. Based on these scores, a top-$k$ operator selects the active indices $\mathcal{I}_{\text{top}k}\subseteq\{1,\dots,r\}$. The adapter weight is then constructed by aggregating the top-$k$ active components:
@@ IV Proposed Zipper-LoRA
To harmonize these objectives, we propose Zipper-LoRA, a family of PEFT modules designed to explicitly allocate and integrate shared and language-specific rank components in a language-aware manner. As illustrated in Fig. 4, all Zipper-LoRA variants leverage a globally shared down-projection matrix AA, ensuring common acoustic-linguistic feature extraction. The fundamental distinction between these variants lies in the rank-wise construction of the up-projection matrix $B$. Intuitively, Zipper-LoRA “zips” together shared and language-specific columns of $B$ to form an effective rank-$r$ adapter for each language. This integration can be realized through either a static allocation or a dynamic rank-wise composition.
To provide a comprehensive understanding of the proposed framework, in the following sections, we first establish the definition of the shared down-projection matrix AA and the rank-wise construction of $B$. We then present the proposed Zipper-LoRA family, ranging from the static language-wise hard split to dynamic binary selection and soft mixing mechanisms. This is followed by the LID-aware contextual routing that manages these rank-wise compositions, concluding with the specialized training strategies tailored for optimizing Zipper-LoRA, such as segment-wise encoding and parameter initialization.
IV-A Definition: Shared AA and Rank-wise Construction of $B$
@@ V-B Configurations
We use Whisper Large-v3 [37] as the speech encoder and Qwen3-1.7B [50] as the decoder-only language model. The modality projector adopts a gated projection module that performs temporal downsampling by a factor of 44. Specifically, it applies two 1D convolution layers (kernel size 44, stride 44) to produce a gate branch and an up-projection branch, fuses them via a SiLU-gated interaction [11], and then applies a linear transformation followed by a residual connection with layer normalization. A final linear projection outputs the speech-conditioned embedding sequence. Combined with the initial $2\times$ downsampling in the speech encoder, this results in an overall $8\times$ temporal reduction, yielding a highly compressed representation at $12.5$ Hz that is fed into the language model.
For parameter-efficient adaptation, we insert low-rank adaptation modules into all linear layers on both the encoder and language-model sides, including the attention projections and feed-forward network layers. We set the rank to $32$ and the scaling factor to $64$ for all methods. For FlyLoRA [57], we activate 88 rank components. For the chunked encoder regime, we randomly sample chunk durations from $\{1,2,4,8\}$ seconds during training and use a fixed chunk duration of 88 seconds during decoding; no overlap is used. The non-chunked regime encodes each utterance with full context. All experiments are trained on 44 NVIDIA A800 GPUs (80 GB) with DeepSpeed ZeRO stage 22 [38], using a learning rate of $2\times 10^{-5}$, a warmup ratio of $10\%$, and a cosine learning-rate schedule. To improve memory efficiency, we pack training samples by constraining the total sequence length to at most 2048 tokens, where the packed length includes the downsampled speech tokens, the text tokens, and reserved prompt tokens; packing is performed within the same language. After packing, we use a batch size of $16$ for chunked training and a batch size of $10$ for non-chunked training.